OpenStudy (anonymous):

can you help me get the formula to solve this. A newspaper sales stand has 200 coins at the end of a day, all dimes and quarters. If the value of the coin is $42.50, how many of each denomination coin were in the stand?

OpenStudy (perl):

Let x = # of dimes , y =# of quarters x+y = 200 .10x + .25y = 42.50

OpenStudy (anonymous):

can you tell me why its .10 and .25

OpenStudy (anonymous):

i came up with my own solution, to my understanding. is it right? 2x + 42.50 = 200 2x = 200-42.50 2x = 157.5 x = 78. 75 i checked it, 2(78.75) + 42.50 = 200 157.5 + 42 .50 = 200 200 = 200